OpenStudy (anonymous):

Factor completely 3x 2 = -5x

OpenStudy (amistre64):

is that 3x^2= -5x?

OpenStudy (amistre64):

x(3x+5) = 0 is complete factored

OpenStudy (anonymous):

yes sorry 3x^2= -5

OpenStudy (amistre64):

3x^2 = -5x 3x^2 +5x = 0 x(3x+5) = 0

OpenStudy (anonymous):

ty